# Efficient Design for Multi-user Downlink Beamforming with Reconfigurable Intelligent Surface

### v2

@@ VII-C Performance Comparison for the MMF Problem
    Relaxed RIS-MMF: Only solve the relaxed problem $\widetilde{{\cal S}}_{o}$ in (VI-A), i.e., Algorithm 2 without the final processing step to obtain the feasible solution ${\bf e}$. We use this method to measure the performance loss in the final processing step in Algorithm 2 and quantify how good the relaxed problem $\widetilde{{\cal S}}_{o}$ is as an approximation of the original problem ${\cal S}_{o}$.
@@
    IP Method: Use MATLAB non-linear optimization solver Fmincon to solve the MMF problem ${\cal S}_{o}$.
  5. 5.
    The alternating projected gradient (APG) method [18]: APG is a recently proposed iterative algorithm to maximize the achievable sum group-rate of all groups. The group rate is defined as the minimum rate among users in the group, which corresponds to the minimum SINR in the group.
Note that the design objective for the APG method [18] is different from the max-min SINR (or rate) objective in our MMF problem. However, for the single-group multicast scenario when $G=1$, the two objectives becomes identical. Since there are no existing RIS-assisted methods designed for the MMF problem, we compare Algorithm 2 with APG in this special single-group case. Besides [18], we note that [17] has also proposed the RIS-assisted multicast beamforming algorithms for the sum-group-rate maximization. Since it is shown in [18] that APG outperforms the method in [17] in both performance and computational complexity, we choose APG as the state-of-the-art method for comparison.
In the following simulation results, we set the BS transmit power $P=10~{}{\text{dBm}}$.
VII-C1 Multiple Groups
Fig. 6 shows the average minimum SINR vs. the number of RIS elements $M$ by different methods for $N=30$. We see that the minimum SINR achieved by Algorithm 2 increases noticeably as $M$ increases, significantly better than random RIS and IP Method, especially when $M>N$. This demonstrates the effectiveness of Algorithm 2 to design RIS beamforming to improve the user channel conditions. Furthermore, we see that Algorithm 2 and Relaxed RIS-MMF provide a nearly identical performance for all values of $M$. This indicates that the solution to $\widetilde{{\cal S}}_{o}$ in Algorithm 2 is mostly feasible to ${\cal S}_{o}$, and the final processing step causes a negligible performance loss. This also demonstrates that the relaxed problem $\widetilde{{\cal S}}_{o}$ is a good approximation of the original problem ${\cal S}_{o}$.
Figure 5: Average minimum SINR vs. $M$ ($N=30$, $(G,K)=(2,2)$).
Figure 5: Average minimum SINR vs. $M$ ($N=30$, $(G,K)=(2,2)$).
Figure 6: Average minimum SINR vs. $N$ ($M=100$, $(G,K)=(2,2)$).
@@ VII-C1 Multiple Groups
Fig. 6 shows the average minimum SINR vs. the number of BS antennas $N$ by different methods for $M=100$. Again, we see that the gain of random RIS and IP Method over no RIS is very small. Our Algorithm 2 provides about $1.5\sim 2$ dB additional SINR gain over these methods for different values of $N$. In Fig. 8, we show the average minimum SINR vs. the number of users $K$ per group for $N=30$ and $M=100$. Algorithm 2 outperforms no RIS, random RIS, and IP Method for all values of $K$, especially for smaller values of $K$. This is expected, since the gain of the optimized multicast beamforming is higher for a smaller user group. The gain reduces as the group size increases, and the difference among algorithms reduces.
Figure 7: Average minimum SINR vs. $K$ ($N=30,M=100$, $(G,K)=(2,2)$).
Figure 7: Average minimum SINR vs. $K$ ($N=30,M=100$, $(G,K)=(2,2)$).
Figure 8: Average minimum SINR vs. $M$ ($N=30$, $(G,K)=(1,4)$).
VII-C2 Single Group
We now compare our proposed algorithm with other algorithms, including the APG method [18], in the single group scenario. Fig. 8 shows the average minimum SINR vs. the number of RIS elements $M$ for $G=1$, $K=4$, and $N=30$. We see that Algorithm 2 outperforms all other algorithms significantly. In particular, our proposed algorithm provides $1\sim 2.5$ dB gain over APG for $M$ ranging from $25$ $\sim 300$.
We also compare the average computation time of different methods in Table II for the setting used in Fig. 8. IP method has the highest computational complexity. Algorithm 2 and APG have a similar computation time as $M$ increases. Although APG is slightly faster than Algorithm 2, the difference is not significant.
TABLE II: MMF – Average Computation Time (sec).
$M$
25
50
100
200
300
APG [18]
0.29
0.71
1.61
2.05
2.7
IP Method
8.73
12.17
36.63
197
476
Algorithm 2
0.45
0.98
2.1
2.6
3.2
VII-C3 Unicast Scenario
We also consider the RIS-assisted downlink beamforming in a unicast scenario with $4$ users by setting $G=4$ and $K=1$. Note that in this case, Algorithm 2 and APG have different design objectives: APG aims to maximize the sum rate of all users, while Algorithm 2 aims to maximize the minimum SINR among users. For comparison, we plot the average minimum SINR and the sum rate performance over $M$ in Figs. 10 and 10, respectively, for $N=30$. We see that Algorithm 2 outperforms the APG method and other methods for both the minimum SINR and the sum rate objectives.
Figure 9: Unicast: Average minimum SINR vs. $M$ ($N=30$, $G=4$ users).
Figure 9: Unicast: Average minimum SINR vs. $M$ ($N=30$, $G=4$ users).
Figure 10: Unicast: Average sum rate vs. $M$ ($N=30$, $G=4$ users).
VIII Conclusion
@@
In this paper, we consider joint BS and RIS beamforming design for RIS-assisted downlink multi-user beamforming, aiming to provide efficient and scalable algorithms. We have shown that the joint BS and RIS beamforming optimization for the QoS problem can be broken down into a BS multicast beamforming QoS problem and an RIS passive multicast beamforming MMF problem, and we have proposed an AMBF approach to solve the two subproblems alternatingly. We have utilized the optimal multicast beamforming structure to solve the BS multicast beamforming QoS problem efficiently. We have also developed a PSA-based fast iterative algorithm to solve the challenging RIS passive multicast MMF problem, which only contains closed-form updates with cheap computation. We have shown that our AMBF approach is also applicable to the unicast scenario, making it an effective method for general RIS-assisted downlink beamforming. For the RIS-assisted MMF problem, based on the relation of the QoS and MMF problems, we have used the optimal BS beamforming structure again to transform the MMF problem and propose a PSA-based iterative algorithm. Similar to that for the QoS problem, The algorithm uses the closed-form updates per iteration, leading to a highly computationally efficient solution. Simulation results show that our proposed algorithms outperform the alternative methods and that the computational complexity is substantially lower than most existing methods and grows linearly with the number of RIS reflective elements and BS antennas.
@@ Appendix A Proof of Proposition 1
Proof:
We can equivalently rewrite ${\cal S}_{o}$ as
@@
For the QoS problem ${\cal P}_{o}$, consider its equivalent formulation ${\cal P}_{2}$ in (III-A) derived in Section III-A. Comparing the above problem and ${\cal P}_{2}$, we see that the objective functions and the constraints in these two problems have the same forms but switched, indicating that they are the inverse problems (and equivalently ${\cal S}_{o}$ and ${\cal P}_{o}$ are the inverse problems). It is straightforward to see that, at the optimality, the respective constraints in ${\cal S}_{o}^{\prime}$ and ${\cal P}_{2}$ are attained with equality. As a result, we have the relations in (27). ∎

### v3

@@ VI-B2 Performance Comparison
    QoS2MMF: Solving ${\cal S}_{o}$ by iteratively solving ${\cal P}_{o}$ as discussed below Proposition 1 in Section V.
@@
    IP Method: Adopt the IP algorithm as a joint optimization method to directly solve the MMF problem ${\cal S}_{o}$ by using the MATLAB non-linear optimization solver Fmincon.
Figure 6: Average minimum SINR vs. $M$ ($N=32$, $(G,K)=(2,3)$).
TABLE II: Average Computation Time (sec) ($N=32$, $(G,K)=(2,3)$).
$M$ 50 100 200 300 400
Algorithm 1 1.19 1.22 1.64 1.75 1.96
QoS2MMF 27.4 32.5 37.8 34.1 58.4
IP Method 7.47 9.84 15.3 22.2 27.5
Note that we select the above methods for comparison as there are no existing methods available for the RIS-assisted MMF problem,555 The sum-group-rate maximization problem is considered as the objective in [27, 28], where the group rate is defined as the minimum rate (i.e., SINR) within the group. Since this objective cannot ensure fairness among groups, the solutions can result in some group having nearly zero rate, which is undesirable. nor a tight upper bound for meaningful comparison. In the following simulation results, we set the BS transmit power $P=10~{}{\text{dBm and $\gamma_{ik}=\gamma=10$~{}dB.}}$..
Fig. 6 shows the average minimum SINR vs. the number of RIS elements $M$ by different methods. We see that the minimum SINR achieved by Algorithm 2 is substantially higher than random RIS and IP method. Also, the SINR gain under Algorithm 2 by increasing $M$ from $50$ to $500$ is nearly 3 dB, significantly more than random RIS and IP method. This shows that Algorithm 2 can provide an effective beamforming solution for RIS elements to improve the performance. Also, the performance of Algorithm 2 and QoS2MMF are close, demonstrating that Algorithm 2 is an effective first-order algorithm that provides a good performance. In contrast, IP method cannot produce an effective solution for the MMF problem and its performance is similar to random RIS. We compare the average computation times of these methods in Table II. The computational advantage of Algorithm 2 over QoS2MMF and IP method can be clearly seen.
@@
Fig. 7 shows the average minimum SINR vs. the number of BS antennas $N$. Since IP method is not effective, we exclude this method in comparison. Again, we see that Algorithm 2 provides about $3$ dB SINR gain over random RIS under $M=100$ for different values of $N$, and its performance is close to that of QoS2MMF. In Fig. 8, we show the average minimum SINR as the number of users $K$ per group increases. Again, Algorithm 2 and QoS2MMF perform similarly. The gain of Algorithm 2 over no RIS or random RIS is consistent for all values of $K$ with $2.5\sim 3$ dB gain.
Figure 7: Average minimum SINR vs. $N$ ($M=100$, $(G,K)=(2,3)$).
Figure 8: Average minimum SINR vs. $K$ ($N=32,M=100$, $G=2$).
VII Conclusion
@@
In this paper, we consider joint BS and RIS beamforming design for RIS-assisted downlink multicasting, aiming to provide efficient and scalable algorithms. We show that the QoS problem can be broken down into a BS multicast beamforming QoS problem and an RIS passive multicast beamforming MMF problem, and we propose an AMBF approach to solve the two subproblems alternatingly. Our proposed fast algorithm for AMBF utilizes the optimal BS mutlicast beamforming structure for computational efficiency and employs a PSA-based fast first-order algorithm to solve the RIS passive multicast MMF problem via iterative closed-form updates. The overall computational complexity grows linearly with the number of BS antennas and RIS elements. For the RIS-assisted MMF problem, we utilize the optimal BS beamforming structure to transform the MMF problem and propose an APSA algorithm using all closed-form updates, leading to a highly computationally efficient solution with complexity linear the number of BS antennas and RIS elements. Simulation results shows that our proposed algorithms provide favorable performance with low cost in computational time as compared with existing methods.
@@ Appendix A Proof of Proposition 1
Proof: We can equivalently rewrite ${\cal S}_{o}$ as
@@
For the QoS problem ${\cal P}_{o}$, consider its equivalent formulation ${\cal P}_{2}$ in (5) derived in Section III-A. Comparing the above problem and ${\cal P}_{2}$, we see that the objective functions and the constraints in these two problems have the same forms but switched, indicating that they are the inverse problems (and equivalently ${\cal S}_{o}$ and ${\cal P}_{o}$ are the inverse problems). It is straightforward to see that, at the optimality, the respective constraints in ${\cal S}_{o}^{\prime}$ and ${\cal P}_{2}$ are attained with equality. As a result, we have the relations in (27).